OpenStudy (precal):

Find the equation of the tangent line to the graph of f(x)=2x+sinx+1 on the interval (0,pi) at the point which is guaranteed by the Mean Value Theorem.

OpenStudy (precal):

\[\frac{ f(\pi)-f(0) }{ \pi-0 }=2\]

OpenStudy (precal):

f ' (c)=2 + cos(c) 2+ cos(c)=2 cos(c)=0 c=pi/2 and (3pi)/2 but pi/2 is the only solution in the interval (0, pi)

OpenStudy (precal):

ok I am stuck at the last part. writing the tangent line

OpenStudy (precal):

ok is it \[y-(\pi+2)=2(x-\frac{ \pi }{ 2 })\]

OpenStudy (precal):

@ganeshie8

OpenStudy (precal):

@agreene

OpenStudy (agreene):

it looks alright to me but I always sucked at making tangent lines lol

OpenStudy (precal):

do you see anything incorrect?

OpenStudy (agreene):

nothing seems wrong.

OpenStudy (precal):

tangent lines are not too bad. you just need the derivative at the given point and you also, need the (x,y) coordinates from that point. so in this case \[f ' (\frac{ \pi }{ 2 })=2\]

OpenStudy (precal):

\[f (\frac{ \pi }{ 2 })=\pi +2\]

OpenStudy (agreene):

yeah, i vaguely remember all of this--I don't use MVT or anything to do with tangent lines in my research and I dont teach calculus 1 so it's a bit hazy lol

OpenStudy (precal):

so we just do simple sub y-y1=m(x-x1) \[y-(\pi +2)=2(x-\frac{ \pi }{ 2 })\]

OpenStudy (ipwnbunnies):

This seems correct to me.

OpenStudy (precal):

Thanks :)